# Robust Decision Transformer: Tackling Data Corruption in Offline RL via Sequence Modeling

### v1

@@ D.1 Impact of Decision Transformer’s Hyperparameters
We aim to understand the robustness characteristic of sequence modeling. The key feature of the DT is its utilization of sequence modeling, with the incorporation of the transformer model. Consequently, we identify two crucial factors that may influence the performance of DT in the context of sequence modeling: the model structure and the input data.
Figure 7: Ablation study on the robustness of DT. (a) Analyzing the influence of block number of DT while keeping sequence length with 20. (b) The effect of sequence length while maintaining 3 transformer blocks. (c) Investigation on how the return-to-go impacts DT. We create the variant DT w/o. Return which excludes the use of return-to-go as inputs.
We explore the influence of the model’s structure by modifying the number of Transformer blocks in DT. As depicted in Figure 7(a), the performance on both MuJoCo and Kitchen tasks consistently improves with an increase in the number of blocks. Therefore, the enhancement can be partially attributed to the increased capabilities of the larger model.
We also examine the impact of the inputs for the sequence model. There are two primary differences between the input of DT and that of conventional BC. DT predicts actions based on historical data and return-to-go elements, whereas BC relies solely on the current states. To assess the influence of these factors, we independently evaluate the impact of the length of history and return-to-go elements. The results in Figure 7(b) and (c) reflect that Kitchen tasks favor longer historical input and are less reliant on return-to-go elements. Conversely, MuJoCo tasks demonstrated a different trend. This discrepancy could be due to the sparse reward structure of Kitchen tasks, where return-to-go elements do not provide sufficient information, leading the agent to rely more heavily on historical data for policy learning. In contrast, MuJoCo tasks, with their denser reward structures, rely more on return-to-go elements for optimal policy learning.
In summary, the robustness of DT can be attributed to the model capacity, history length, and the return-to-go conditioning. Based on this findings on the robustness of DT, we adopt a sequence length of 20 and a block number of 3 as the default implementation for all sequence modeling methods, including DT, DeFog, and RDT.
@@ D.2 Ablation Study on Reward Prediction
The nature of the sequential model allows us to predict not only actions but also other elements such as state and return-to-go. Previous studies have shown that predicting these additional elements does not significantly enhance performance Chen et al., (2021). However, we find that predicting rewards can provide advantages in the context of data corruption. We introduce DT(RP), a variant of the original Decision Transformer (DT), which predicts rewards in addition to actions. As illustrated in Figure 8, predicting rewards indeed improves performance under state and action corruption conditions. However, DT(RP) performs poorly under reward corruption due to the degraded quality of reward labels. This issue can be mitigated via embedding dropout, Gaussian weighted learning, and iterative data correction in RDT.
@@
D.3 Ablation Study on Hyperparameters
We conducted experiments to investigate the impact of hyperparameters of the Robust Decision Transformer (RDT) under scenarios of action corruption.
In our study of the embedding dropout technique, depicted in Figure 9(a), we discovered that applying a moderate dropout probability (e.g., 0.1) to embeddings enhances resilience against data corruption. However, a larger dropout rate ($\geq$ 0.5) can lead to performance degradation.
As shown in Figure 9(b), regarding the Gaussian weighted coefficient $\beta_{a}$, the walker2d task favors lower values of $\beta_{a}$ such as 0.1, while the kitchen and relocate tasks favor larger $\beta_{a}$ values, suggesting that these tasks are more impacted by Gaussian weight learning.
In terms of the iterative data correction threshold $\zeta$, the results are highlighted in Figure 9(c). The walker2d and kitchen tasks favor lower values of $\zeta$, whereas the relocate task exhibits the opposite preference. Notably, in the relocate task, RDT shows markedly enhanced performance after correcting outliers (with $\zeta\geq 5$) in the dataset.
(a) Embedding drop $p$
(b) Gaussian weight coefficient $\beta_{a}$
(c) Iterative correction threshold $\zeta$
@@ D.3 Ablation Study on Hyperparameters
D.4 Additional Results under Observation Perturbation during the Testing Phase
@@
To further demonstrate the robustness of RDT, we conducted additional evaluations under observation perturbation during the testing phase. Specifically, we compared different algorithms trained with offline datasets under conditions of state or reward corruption in perturbed environments. As shown in Figures 10 and 11, RDT consistently maintained superior performance and robustness across observation perturbations in both the Kitchen and Adroit tasks.
(a) "Random" observation perturbation.
(b) "Action diff” observation perturbation.
@@
(a) "Random" observation perturbation.
(b) "Action diff" observation perturbation.
@@ D.5 Training Time
We trained all algorithms on the same P40 GPU, and the training times are presented in Table 4. As expected, BC has the shortest epoch time due to its simplicity. Notably, RDT requires only slightly more time per epoch than DT because it incorporates three robust techniques. Although RIQL requires less time per epoch than RDT, it necessitates 10 times more epochs to converge and still does not achieve satisfactory performance in the limited data regime. In conclusion, RDT achieves superior performance and robustness without imposing significant computational costs in terms of total training time.
Table 4: Training time on the "walker2d-medium-replay" dataset.
Method BC CQL UWMSG RIQL DT RDT (ours)
Epoch Num 1000 1000 1000 1000 100 100
Epoch Time (s) 3.7 33.8 22.2 14.9 44.0 46.1
Total Time (h) 1.03 9.39 6.17 4.14 1.22 1.28
D.6 Detailed Results for the Main Experiments
The detailed outcomes of each task, under both random and adversarial data corruption, can be found in Table 5 and Table 6 respectively. For HalfCheetah, Hopper, and Walker2D tasks, we have utilized the "medium-replay-v2" datasets, while the "expert-v0" datasets have been used for Door, Hammer, and Relocate tasks. Notably, RDT consistently outperforms baselines in the majority of tasks, achieving the best average performance and demonstrating its robustness and superiority, particularly under the adversarial data corruption.

### v3

@@ Appendix D Additional Experiments
To provide a comprehensive understanding of our work, we conduct several additional analyses. First, we investigate the key hyperparameters that influence the robustness of DT. Next, we discuss the impact of reward prediction on DT. Following this, we demonstrate how these hyperparameters affect RDT and assess the robustness of RDT under various scales of data corruption. We then compare the training time between RDT and baseline models. Finally, we present detailed results of RDT under both training-time corruption and testing-time perturbation.
@@ D.1 Impact of Decision Transformer’s Hyperparameters
We aim to understand the robustness characteristic of sequence modeling. The key feature of the DT is its utilization of sequence modeling, with the incorporation of the Transformer model. Therefore, we identify two crucial factors that may influence the performance of DT in the context of sequence modeling: the model structure and the input data.
Figure 8: Ablation study on the robustness of DT. (a) Analyzing the influence of Transformer block number while keeping sequence length with 20. (b) The effect of sequence length while maintaining 3 Transformer blocks. (c) Investigation on how the return-to-go impacts DT. We create the variant DT w/o. Return, which excludes the use of return-to-go as inputs.
We explore the influence of the model’s structure by modifying the number of Transformer blocks in DT. As depicted in Figure 8(a), the performance on both MuJoCo and Kitchen tasks consistently improves with increased number of blocks. Therefore, the enhancement can be partially attributed to the increased capabilities of the larger model.
We also examine the impact of the inputs for the sequence model. There are two primary differences between the input of DT and that of conventional BC. DT predicts actions based on historical data and return-to-go elements, whereas BC relies solely on the current states. To assess the influence of these factors, we independently evaluate the impact of the length of history and return-to-go elements. The results in Figure 8(b) and (c) reflect that Kitchen tasks favour longer historical input and are less reliant on return-to-go elements. Conversely, MuJoCo tasks demonstrated a different trend. This discrepancy could be due to the sparse reward structure of Kitchen tasks, where return-to-go elements do not provide sufficient information, leading the agent to rely more heavily on historical data for policy learning. In contrast, MuJoCo tasks, with their denser reward structures, rely more on return-to-go elements for optimal policy learning.
In summary, the robustness of DT can be attributed to the model capacity, history length, and return-to-go conditioning. Based on these findings on the robustness of DT, we adopt a sequence length of 20 and a block number of 3 as the default implementation for all sequence modeling methods, including DT, DeFog, and RDT.
@@ D.2 Ablation Study on Reward Prediction
The nature of the sequence model allows us to predict not only actions but also other elements, such as state and return-to-go. Previous studies have shown that predicting these additional elements does not significantly enhance performance Chen et al. (2021). However, we find that predicting rewards can provide advantages in the context of data corruption. We introduce DT(RP), a variant of the original DT, which predicts rewards in addition to actions. As illustrated in Figure 9, predicting rewards indeed improves performance under random state and action corruption conditions. However, DT(RP) performs poorly under reward corruption due to the degraded quality of reward labels. This issue can be mitigated via embedding dropout, Gaussian weighted learning, and iterative data correction in RDT.
@@
D.3 Ablation Study on Hyperparameters of RDT
We conduct experiments to investigate the impact on hyperparameters of RDT under scenarios of random action corruption.
In the study of the embedding dropout technique, depicted in Figure 10(a), we discover that applying a moderate dropout probability (e.g., 0.1) to embeddings enhances resilience against data corruption. However, a larger dropout rate ($\geq$ 0.5) can lead to performance degradation.
As shown in Figure 10(b), regarding the Gaussian weighted coefficient $\beta_{a}$, the walker2d task favours lower values of $\beta_{a}$ such as 0.1, while the kitchen-complete and relocate tasks favour larger $\beta_{a}$ values, suggesting that these tasks are less impacted by Gaussian weight learning.
In terms of the iterative data correction threshold $\zeta$, the results are highlighted inFigure 10(c). The walker2d and kitchen-complete tasks favour lower values of $\zeta$, whereas the relocate task exhibits the opposite preference. Notably, in the relocate task, RDT shows markedly enhanced performance after correcting outliers (with $\zeta\geq 5$) in the dataset.
(a) Embedding drop $p$.
(b) Gaussian weight coefficient $\beta_{a}$.
(c) Iterative correction threshold $\zeta$.
@@ D.3 Ablation Study on Hyperparameters of RDT
D.4 Varying Corruption Rates and Scales
We have evaluated the effectiveness of RDT under a data corruption rate of $0.3$ and a scale of $1.0$ in Section 4.2. We further examine the robustness of RDT under different corruption rates from $\{0.0,0.1,0.3,0.5\}$ and scales from $\{0.0,1.0,2.0\}$. As illustrated in Figure 11, RDT consistently delivers superior performance compared to other baselines across different corruption rates and scales. However, we also observed a decline in RDT’s performance when it encountered state corruption with high rates. This decline can be attributed to the distortion of the state’s distribution caused by the high corruption rates, resulting in a significant deviation from the clean test environment to the corrupted datasets. Enhancing the robustness of RDT against state corruption, especially under high corruption rates and scales, holds potential for future advancements.
Figure 11: Results under various corruption rates (a) and scales (b) on the “walker2d” task.
D.5 Detailed Results for the Experiments on Adversarial Attack
The detailed outcomes of each task under adversarial data corruption are presented in Table 5. Notably, RDT outperforms baseline models in the majority of tasks, achieving the best average performance. In particular, RDT consistently outperforms DT, demonstrating the effectiveness of our three robust techniques.
Table 5: Detailed comparative results under adversarial data corruption.
Attack Task BC RBC DeFog CQL UWMSG RIQL DT RDT State halfcheetah (10%) 2.1$\pm$ 0.6 2.4$\pm$ 0.3 4.8$\pm$ 0.7 10.0$\pm$ 1.6 7.2$\pm$ 2.6 3.6$\pm$ 0.5 5.4$\pm$ 0.7 7.5$\pm$ 0.9 hopper (10%) 19.1$\pm$ 3.2 18.7$\pm$ 4.4 21.8$\pm$ 7.4 1.8$\pm$ 0.8 14.7$\pm$ 3.4 19.1$\pm$ 6.9 38.9$\pm$ 2.9 39.1$\pm$ 4.6 walker2d (10%) 8.7$\pm$ 1.8 7.1$\pm$ 0.2 12.0$\pm$ 3.6 0.7$\pm$ 1.5 5.1$\pm$ 2.3 10.4$\pm$ 1.9 21.3$\pm$ 1.5 24.2$\pm$ 3.2 kitchen-complete 18.9$\pm$ 7.0 20.8$\pm$ 9.9 43.5$\pm$ 5.5 4.6$\pm$ 4.8 0.0$\pm$ 0.0 51.9$\pm$ 3.3 37.6$\pm$ 3.2 61.1$\pm$ 2.0 kitchen-partial 27.1$\pm$ 4.2 30.8$\pm$ 3.8 8.9$\pm$ 5.2 0.0$\pm$ 0.0 0.0$\pm$ 0.0 35.4$\pm$ 5.8 37.9$\pm$ 6.2 41.6$\pm$ 5.1 kitchen-mixed 24.1$\pm$ 3.1 35.2$\pm$ 3.7 7.6$\pm$ 6.3 2.0$\pm$ 3.5 0.0$\pm$ 0.0 33.9$\pm$ 11.2 38.1$\pm$ 6.0 44.5$\pm$ 5.1 door (1%) 71.1$\pm$ 8.8 66.1$\pm$ 10.2 102.2$\pm$ 1.4 -0.3$\pm$ 0.0 -0.3$\pm$ 0.1 47.3$\pm$ 24.8 98.2$\pm$ 4.8 104.0$\pm$ 1.0 hammer (1%) 82.5$\pm$ 11.1 79.3$\pm$ 8.4 102.7$\pm$ 8.8 0.2$\pm$ 0.0 0.1$\pm$ 0.1 69.1$\pm$ 23.4 90.1$\pm$ 11.9 110.9$\pm$ 9.0 relocate (1%) 27.0$\pm$ 2.6 15.5$\pm$ 7.2 22.1$\pm$ 3.5 -0.3$\pm$ 0.0 -0.3$\pm$ 0.0 17.1$\pm$ 9.5 67.6$\pm$ 2.6 71.3$\pm$ 2.5 Average 31.2 30.6 36.2 2.1 2.9 32.0 48.4 56.0 Action halfcheetah (10%) -0.1$\pm$ 0.1 -0.0$\pm$ 0.2 8.9$\pm$ 1.0 19.0$\pm$ 2.2 6.7$\pm$ 1.1 0.8$\pm$ 0.4 2.4$\pm$ 0.1 17.1$\pm$ 3.4 hopper (10%) 10.9$\pm$ 3.7 11.1$\pm$ 2.2 16.1$\pm$ 10.4 1.5$\pm$ 0.5 37.4$\pm$ 4.0 17.1$\pm$ 1.4 28.8$\pm$ 3.3 31.6$\pm$ 1.8 walker2d (10%) 1.9$\pm$ 0.7 1.9$\pm$ 0.5 5.8$\pm$ 2.4 1.4$\pm$ 1.0 6.7$\pm$ 1.0 3.3$\pm$ 1.2 9.0$\pm$ 1.4 16.0$\pm$ 1.7 kitchen-complete 6.6$\pm$ 0.6 5.6$\pm$ 3.0 11.5$\pm$ 11.1 1.8$\pm$ 1.9 0.0$\pm$ 0.0 20.6$\pm$ 4.5 11.5$\pm$ 2.3 25.5$\pm$ 1.7 kitchen-partial 5.8$\pm$ 3.6 6.9$\pm$ 2.3 0.0$\pm$ 0.0 1.0$\pm$ 1.7 0.0$\pm$ 0.0 0.0$\pm$ 0.0 1.9$\pm$ 1.4 30.8$\pm$ 4.5 kitchen-mixed 6.4$\pm$ 7.4 22.1$\pm$ 4.3 0.1$\pm$ 0.2 0.0$\pm$ 0.0 0.0$\pm$ 0.0 3.4$\pm$ 4.5 2.8$\pm$ 3.4 45.9$\pm$ 4.6 door (1%) 6.0$\pm$ 1.3 11.8$\pm$ 5.7 102.8$\pm$ 1.1 -0.3$\pm$ 0.1 -0.2$\pm$ 0.1 38.7$\pm$ 21.0 62.2$\pm$ 7.4 97.6$\pm$ 5.5 hammer (1%) 16.3$\pm$ 10.2 44.0$\pm$ 12.2 37.9$\pm$ 9.9 0.2$\pm$ 0.0 0.2$\pm$ 0.1 88.1$\pm$ 14.6 62.2$\pm$ 16.7 107.7$\pm$ 1.3 relocate (1%) 2.7$\pm$ 2.7 6.2$\pm$ 3.4 14.7$\pm$ 4.7 -0.2$\pm$ 0.1 -0.3$\pm$ 0.0 0.4$\pm$ 0.5 18.0$\pm$ 6.0 35.7$\pm$ 3.1 Average 6.3 12.2 22.0 2.7 5.6 19.2 22.1 45.3 Reward halfcheetah (10%) 2.4$\pm$ 0.2 2.9$\pm$ 1.0 15.0$\pm$ 2.2 31.2$\pm$ 1.6 1.9$\pm$ 0.2 10.9$\pm$ 1.6 9.5$\pm$ 0.6 23.3$\pm$ 5.5 hopper (10%) 19.7$\pm$ 2.8 19.3$\pm$ 3.1 15.6$\pm$ 8.4 1.8$\pm$ 0.0 28.6$\pm$ 9.9 34.5$\pm$ 5.1 37.2$\pm$ 6.5 37.2$\pm$ 6.6 walker2d (10%) 9.7$\pm$ 1.5 8.3$\pm$ 1.7 4.0$\pm$ 1.9 2.2$\pm$ 2.4 9.7$\pm$ 3.1 9.3$\pm$ 1.2 29.0$\pm$ 3.6 35.2$\pm$ 2.9 kitchen-complete 36.0$\pm$ 11.5 38.2$\pm$ 5.0 48.0$\pm$ 2.4 3.4$\pm$ 5.8 0.0$\pm$ 0.0 51.5$\pm$ 2.6 45.0$\pm$ 5.0 65.9$\pm$ 2.0 kitchen-partial 34.1$\pm$ 1.4 39.1$\pm$ 1.8 9.5$\pm$ 7.9 0.0$\pm$ 0.0 0.0$\pm$ 0.0 37.2$\pm$ 8.5 45.1$\pm$ 6.1 46.1$\pm$ 1.6 kitchen-mixed 38.9$\pm$ 1.4 47.1$\pm$ 2.0 3.5$\pm$ 3.3 0.0$\pm$ 0.0 0.0$\pm$ 0.0 48.9$\pm$ 4.7 54.1$\pm$ 2.3 56.1$\pm$ 2.8 door (1%) 76.0$\pm$ 5.9 75.0$\pm$ 9.0 102.2$\pm$ 1.0 -0.3$\pm$ 0.1 -0.2$\pm$ 0.1 73.4$\pm$ 11.3 98.9$\pm$ 2.8 101.4$\pm$ 0.6 hammer (1%) 97.1$\pm$ 8.3 99.0$\pm$ 11.8 92.1$\pm$ 14.0 0.2$\pm$ 0.0 0.1$\pm$ 0.1 69.1$\pm$ 27.7 97.1$\pm$ 11.3 110.9$\pm$ 1.6 relocate (1%) 36.1$\pm$ 8.6 32.2$\pm$ 3.2 48.1$\pm$ 5.0 -0.3$\pm$ 0.1 -0.3$\pm$ 0.0 18.1$\pm$ 6.0 76.8$\pm$ 9.1 77.2$\pm$ 4.2 Average 38.9 40.1 37.6 4.3 4.4 39.2 54.7 61.5 Average over all tasks 25.5 27.7 31.9 3.0 4.3 30.1 41.7 54.3
D.6 Comparison Results on MuJoCo Tasks with Full Dataset
We further evaluate the robustness of RDT on MuJoCo tasks using the full dataset. As shown in Table 6, RIQL is comparable to DT in this setting, while RDT achieves the highest overall average score. These findings further demonstrate the effectiveness of our proposed robust techniques across different dataset scales.
Table 6: Results under random data corruption on MuJoCo tasks with 100% dataset.
Attack Task BC RBC DeFog CQL UWMSG RIQL DT RDT State halfcheetah 32.4$\pm$ 1.9 33.0$\pm$ 0.9 14.9$\pm$ 7.1 23.7$\pm$ 4.4 2.2$\pm$ 0.6 19.9$\pm$ 2.1 27.5$\pm$ 2.5 30.8$\pm$ 1.8 hopper 27.1$\pm$ 8.6 26.6$\pm$ 2.9 21.1$\pm$ 6.8 34.4$\pm$ 22.8 21.6$\pm$ 8.0 34.0$\pm$ 13.4 51.3$\pm$ 14.0 56.6$\pm$ 2.9 walker2d 19.2$\pm$ 4.7 16.8$\pm$ 6.0 17.4$\pm$ 2.4 33.7$\pm$ 10.8 0.8$\pm$ 1.7 14.2$\pm$ 1.2 47.6$\pm$ 4.9 53.4$\pm$ 4.0 Average 26.2 25.4 17.8 30.6 8.2 22.7 42.1 46.9 Action halfcheetah 34.2$\pm$ 1.4 34.6$\pm$ 1.9 34.2$\pm$ 1.8 43.9$\pm$ 0.4 49.9$\pm$ 1.2 41.3$\pm$ 1.4 35.1$\pm$ 2.4 37.4$\pm$ 1.4 hopper 22.5$\pm$ 4.8 28.8$\pm$ 10.1 40.8$\pm$ 11.6 29.8$\pm$ 8.8 39.0$\pm$ 6.8 59.6$\pm$ 12.7 63.2$\pm$ 5.6 63.2$\pm$ 5.8 walker2d 18.3$\pm$ 2.6 21.3$\pm$ 3.6 50.7$\pm$ 6.0 1.8$\pm$ 3.6 53.3$\pm$ 16.8 72.6$\pm$ 17.3 59.0$\pm$ 5.5 61.4$\pm$ 1.7 Average 25.0 28.2 41.9 25.2 47.4 57.8 52.4 54.0 Reward halfcheetah 35.9$\pm$ 0.5 34.7$\pm$ 1.4 32.6$\pm$ 2.3 43.3$\pm$ 0.3 37.2$\pm$ 5.8 42.7$\pm$ 0.9 38.4$\pm$ 0.3 39.6$\pm$ 0.8 hopper 26.6$\pm$ 6.9 29.5$\pm$ 8.6 29.2$\pm$ 18.3 29.9$\pm$ 7.2 65.4$\pm$ 26.0 59.6$\pm$ 4.8 54.3$\pm$ 10.7 70.9$\pm$ 8.7 walker2d 23.5$\pm$ 9.8 22.9$\pm$ 6.9 40.1$\pm$ 9.4 38.3$\pm$ 25.9 51.3$\pm$ 10.6 79.0$\pm$ 5.8 64.2$\pm$ 4.0 66.7$\pm$ 4.0 Average 28.7 29.0 34.0 37.2 51.3 60.4 52.3 59.0 Average over all tasks 26.6 27.6 31.2 31.0 35.6 47.0 48.9 53.3
Our reported results for RIQL on MuJoCo tasks using the full 100% dataset show some discrepancies compared to the original paper (Yang et al., 2024b). To investigate this, we plotted the learning curve of RIQL under an action attack, as shown in Figure 12. In comparison to Figure 17 in the RIQL paper, we found that RIQL may require more training epochs to converge when using the full dataset, which is reasonable. Given our focus on the limited data setting (10% for MuJoCo), we observed that 1,000 epochs are sufficient for both RDT and baselines like RIQL to converge, as detailed in Section D.8. Therefore, we use 1,000 training epochs by default, which, however, may not be optimal for the full dataset setting.
Figure 12: Learning curve of RIQL under action attack.
@@ D.6 Comparison Results on MuJoCo Tasks with Full Dataset
D.7 Additional Results under Observation Perturbation during Testing Phase
@@
To further demonstrate the robustness of RDT, we conduct additional evaluations under observation perturbations during the testing phase. Specifically, we compare different algorithms trained with offline datasets under conditions of random state, action, or reward corruption in environments with observation perturbations. As shown in Figures 13, 14 and 15, RDT consistently maintains superior performance and robustness across observation perturbations in both the Kitchen and Adroit tasks. We leave the investigation of robustness improvements over MuJoCo tasks for future work.
(a) Random observation perturbation.
(b) Action Diff observation perturbation.
@@
(a) Random observation perturbation.
(b) Action Diff observation perturbation.
Figure 14: Performance under varying observation perturbation scales during the testing phase. All the algorithms are trained under random action corruption during the training phase.
(a) Random observation perturbation.
(b) Action Diff observation perturbation.
@@ D.8 Training Time
We train all algorithms on the same P40 GPU, and the training times are presented in Table 7. As expected, BC has the shortest epoch time due to its simplicity. Notably, RDT requires only slightly more time per epoch than DT because it incorporates three robust techniques. Although RIQL requires less time per epoch than RDT, it necessitates 10 times more epochs to converge and still does not achieve satisfactory performance in the limited data regime. In conclusion, RDT achieves superior performance and robustness without imposing significant computational costs in terms of total training time.
Table 7: Training time on the “walker2d-medium-replay-v2” dataset.
Method BC CQL UWMSG RIQL DT RDT Epoch Num 1000 1000 1000 1000 100 100 Epoch Time (s) 3.7 33.8 22.2 14.9 44.0 46.1 Total Time (h) 1.03 9.39 6.17 4.14 1.22 1.28
D.9 Additional Ablation Study on Three Robust Techniques
We investigate the impact of each individual technique in Section 4.4 and demonstrate that integrating all techniques is essential for achieving optimal robustness. To provide a more comprehensive analysis of our proposed techniques, we create the variants RDT w/o ED, RDT w/o GWL, and RDT w/o IDC, which eliminate the components Embedding Dropout, Gaussian Weighted Learning, and Iterative Data Correction from RDT, respectively. As illustrated in Figure 16, the results are consistent with the findings in Section 4.4. This consistency reinforces our initial conclusion that the integration of all proposed techniques is crucial for achieving optimal robustness.
Figure 16: Additional ablation study on the impact of proposed techniques.
D.10 Evaluation under Various Dataset Scales
We conducted experiments across various dataset scales to understand the impact of dataset scales. Specifically, we evaluate RDT on the MuJoCo tasks with 20% and 50% of the “medium-replay-v2” dataset and the Adroit tasks with 5% and 10% of the “expert-v0” dataset. The comparison results are shown in Figure 17. Our empirical findings indicate that: (1) The overall performance of all methods improves as the dataset size increases. (2) Moreover, temporal-difference-based methods like RIQL are comparable to vanilla sequence modeling methods like DT when the dataset is large, but they are less preferable when the dataset size is limited. (3) Notably, RDT consistently outperforms other baselines across all dataset sizes, validating the effectiveness and importance of our proposed robustness techniques.
(a) Results on MuJoCo tasks.
(b) Results on Adroit tasks.
Figure 17: Performance of different algorithms across various dataset scales.
D.11 Performance under MuJoCo Dataset with Different Quality Levels
To further validate the robustness of the RDT across varying dataset quality levels, we conduct evaluations using both the “expert-v2” and “medium-v2” datasets on MuJoCo tasks. We randomly sample $2\%$ for both datasets, respectively, resulting in both datasets having a similar size of approximately $2\times 10^{4}$ transitions. We include RIQL, DeFog, and DT as baselines, and the results are presented in Tables 8 and 9. Notably, RDT consistently outperforms the other baselines across datasets of different quality levels. This superior performance reinforces RDT’s capability to efficiently handle a range of dataset qualities.